# MMEarth: Exploring Multi-Modal Pretext Tasks For Geospatial Representation Learning

### v1

@@ 6.3 Limitations
While this study introduces a new dataset and methodology to advance the use of multiple modalities for representation learning, some limitations exist. First, we rely on GEO-Bench for downstream evaluation with five datasets covering a range of tasks and geographic regions. Further evaluation on diverse application domains from various geographical regions could reveal further insights in which scenarios our proposed multi-modal pretext tasks are beneficial. Second, we run our pretraining for 200 epochs due to restricted compute resources. For the same reason we only train encoders with limited number of parameters. Prior work exploring MAEs for representation learning pretrain for 800 or 1600 epochs. Pretraining for 200 epochs could be suboptimal for larger encoders. Third, our MP-MAE is not explicitly trained to be robust to varying number of input channels. We have trained the MP-MAE with multi-spectral Sentinel-2 images (12 bands). It needs to be investigated how our models transfer to downstream tasks that use RGB channels only. While this seems to be an artefact of benchmark datasets, rather than a real-world scenario when working with Sentinel-2 images, prior work proposed input masking strategies to make a model robust against missing channels [45], which could be built into our framework.
@@ 7 Conclusion
Prior work has mostly focused on exploiting geolocation and time for representation learning with EO data. We move a step forward and present MMEarth, a global multi-modal pretraining dataset for geospatial representation learning. With 12 modalities sampled from 1.2M locations, we propose a Multi-Pretext Masked Autoencoder (MP-MAE) approach based on ConvNeXt V2 to study the potential of aligned multi-modal data to learn better representations for Sentinel-2 optical satellite images. First, we show how to adapt the ConvNeXt V2 encoder to improve domain specific pretraining with medium-resolution satellite images. Second, we demonstrate that multi-modal pretext tasks are beneficial to learn better representations for Sentinel-2. Both fine-tuning and linear probing is improved, the latter profits substantially. This is promising for real-world applications with limited training data. However, further research is needed to close the gap between fine-tuning and linear-probing performance.
@@
We greatly appreciate the open data policies of the Copernicus program and its partners ESA and ECMWF. We thank Google Earth Engine for hosting the data and providing free access. VN, CI, SB, and NL acknowledge support by the Pioneer Centre for AI, Danish National Research Foundation grant number P1. SO, AK, CI, and NL acknowledge support by the research grant DeReEco (grant number 34306) from Villum Foundation.
@@ 0.A.1 Sampling strategy
ERA5. We selected 4 bands from the ERA5 dataset on GEE (average temperature, minimum temperature, maximum temperature and total precipitation). These statistics are monthly. Additionally, we computed the yearly average, minimum, maximum, and total, respectively, for each of these statistics. We collected 3 sets of statistics for the same month, previous month, and full year corresponding to the Sentinel-2 observation date ($3\ sets\times 4\ bands=12\ bands$).
Geolocation. We use a cyclic encoding of the latitude and longitude corresponding to the center point as follows:
$Lat_{sin}=\sin{(2\pi Lat/360)}$ (3)
$Lat_{cos}=\cos{(2\pi Lat/360)}$ (4)
$Lon_{sin}=\sin{(2\pi Lon/360)}$ (5)
$Lon_{cos}=\cos{(2\pi Lon/360)}$ (6)
@@
$Month_{sin}=\sin{(2\pi m/12)}$ (7)
$Month_{cos}=\cos{(2\pi m/12)}$ (8)
where $m$, corresponds to the month as an integer from 1 to 12.
@@ 0.A.2 Licenses of data sources
Appendix 0.B Implementation details for evaluation
So2Sat20k only provides 10 instead of 12 bands. However, to make use of these additional multi-spectral bands, we filled in the missing bands (B1 and B9) with a copy of the band with the closest wavelength. I.e., we use a copy of B2 for B1, and B8A for B9. This could lead to a lower linear probing performance but should not affect results much when fine-tuning the entire encoder. To adapt to these slight distribution shifts in band B1 and B9, we ran an extended linear probing, where the parameters of the first convolution layer are updated in addition to the last linear layer. Results from Table 2(b) show that this extended linear probing improved performance compared to linear probing on So2Sat20k.
Table A2: Extended linear probing on So2Sat20k. Extended linear probing experiments with updating the parameters of the first convolution layer in addition to the last layer.
(a) Pretraining on MMEarth-S2
So2Sat20k
LP 28.6
LP + first layer 37.0
(b) Pretraining on MMEarth
So2Sat20k
LP 44.2
LP + first layer 49.8
@@ Appendix 0.C Taster pretraining datasets: MMEarth100k and MMEarth64
To facilitate research in self-supervised learning with limited compute resources, we provide two subsets of the MMEarth dataset: MMEarth100k and MMEarth64. The MMEarth100k is a random subset of 100k locations with the full 128$\times$128 pixels and has a data volume of roughly 48GB. The MMEarth64 contains all 1.2M locations, but all modalities are center cropped to obtain rasters of 64$\times$64 pixels, which reduces the data volume roughly by factor 4 from 639GB to 163GB.
We provide results using these taster subsets to pretrain our MP-MAE approach with the same setting as used for pretraining on the full MMEarth dataset in Table A3. The only change required for MMEarth64 is the patch size. As the image size is halved from 112 to 56 pixels, we also half the patch size from 16 to 8 pixels which keeps the number of patches the same. We observe that pretraining on MMEarth64 performs comparably to MMEarth, and even slightly better in some tasks. Pretraining on MMEarth100k lead to a substantial drop in performance, but still outperforms the ImageNet pretrained models on BigEarth20k and SAcrop3k. From these initial results, we conclude that the diversity and number of samples is more important than the size of the receptive field for learning generalizing representations.
Table A3: MP-MAE results on taster datasets.
Pretrain data Image size Patch size BigEarth20k FT/LP So2Sat20k FT/LP Cashew1k FT SAcrop3k FT ImageNet 224 32 72.9/57.2 44.5/34.9 77.5 27.4 MMEarth 112 16 80.0/68.2 55.9/44.2 81.4 34.5 MMEarth100k 112 16 75.5/63.3 42.9/31.4 68.4 30.6 MMEarth64 56 8 82.8/67.4 56.9/42.3 81.6 35.8
@@ 0.D.1.1 Analysis of the estimated task uncertainty
0.D.2 Input normalization for downstream tasks
During pretraining, we compute the normalization statistics across the whole dataset for each modality (global statistics). In the case of Sentinel-2, we additionally compute separate statistics for L1C and L2A data. When evaluating on downstream tasks, we apply the normalization statistics computed over the downstream dataset. This can introduce a distribution shift when performing FT/LP. To avoid this, we run an experiment by applying the same global stats to the downstream evaluation tasks. Specifically, if the downstream task has been collected with Sentinel-2 L2A data (e.g. BigEarth20k), we apply the global statistics for L2A on this task. Results from Table LABEL:tab:input_norm indicate that in the case of BigEarth20k, using global pretraining statistics can help improve the performance, but the improvement is not consistent for the other downstream tasks.
Table A4: Impact of input normalization for downstream learning. Results for the Atto model. Domain specific pretraining allows to transfer the input statistics of the pretraining to the downstream tasks. For downstream tasks using L2A (or L1C) data, we apply the corresponding pretraining L2A (or L1C) statistics. This reduces domain shifts between the pretraining and downstream tasks and facilitates FT and LP. We omit SAcrop3k from these experiments because the benchmark dataset contains preprocessed data that are neither L1A nor L1C reflectance values.
Pretrain FT/LP input normalization BigEarth20k FT/LP So2Sat20k FT/LP Cashew1k FT MMEarth FT/LP train stats 80.0/68.2 55.9/44.2 81.4 MMEarth Pretraining stats L2A 82.3/71.3 - 79.3 MMEarth Pretraining stats L1C - 55.6/43.4 -
Table A5: Loss aggregation strategy. We evaluate the effect of the two loss aggregation strategies: Equal weighting vs. task uncertainty weighting.
Pretrain Model BigEarth20k FT/LP So2Sat20k FT/LP Cashew1k FT SAcrop3k FT MP-MAE (ours) Atto-Uncertainty 80.0/68.2 55.9/44.2 81.4 34.5 MP-MAE (ours) Atto-Equal 80.9/69.3 58.8/41.5 80.1 34.2
0.D.3 Qualitative reconstruction results

### v2

@@ 6.3 Limitations
While this study introduces a new dataset and methodology to advance the use of multiple modalities for representation learning, some limitations exist. First, we rely on GEO-Bench for downstream evaluation with five datasets covering a range of tasks and geographic regions. Further evaluation on diverse application domains from various geographical regions could reveal insights about which scenarios our proposed multi-modal pretext tasks are beneficial in at a global scale. Second, we run our pretraining for 200 epochs due to restricted compute resources. For the same reason we only train encoders with a limited number of parameters. Prior work exploring MAEs for representation learning pretrains for 800 or 1600 epochs. Third, our MP-MAE is not explicitly trained to be robust to varying numbers of input channels. We have trained the MP-MAE with multi-spectral Sentinel-2 images (12 bands). It needs to be investigated how our models transfer to downstream tasks that use e.g. RGB only. While this seems to be an artefact of benchmark datasets, rather than a real-world scenario when working with Sentinel-2 images, prior work proposed input masking strategies to make a model robust against missing channels [46], which could be built into our framework. Fourth, we are not making use of the validation dataset for early stopping during the FT/LP evaluation, which could improve results overall. Lastly, we only reported single trials. Running the FT/LP experiments multiple times would allow us to average out random fluctuations.
@@ 7 Conclusion
Prior work has mostly focused on exploiting geolocation and time for representation learning with EO data. We move a step forward and present MMEarth, a global multi-modal pretraining dataset for geospatial representation learning. With 12 modalities sampled from 1.2M locations, we propose a Multi-Pretext Masked Autoencoder (MP-MAE) approach based on ConvNeXt V2 to study the potential of aligned multi-modal data to learn better representations for Sentinel-2 optical satellite images. First, we show how to adapt the ConvNeXt V2 encoder to improve domain-specific pretraining with medium-resolution satellite images. Second, we demonstrate that multi-modal pretext tasks are beneficial for learning better representations for Sentinel-2. Both fine-tuning and linear probing are improved; the latter profits substantially in the few-shot setting. This is promising for real-world applications with limited training data. However, further research is needed to close the gap between the fine-tuning and linear probing performance.
@@
We thank Lucia Gordon for the valuable feedback. We greatly appreciate the open data policies of the Copernicus program and its partners ESA and ECMWF. We thank Google Earth Engine for hosting the data and providing free access. This work was supported in part by the Pioneer Centre for AI, DNRF grant number P1. The authors AK, CI, and NL acknowledge support by the research grant DeReEco (grant number 34306) from Villum Foundation. SO and CI acknowledge support by the research grant Global Wetland Center (grant number NNF23OC0081089) from Novo Nordisk Foundation. CI and SB acknowledge support by the European Union project ELIAS (grant agreement number 101120237). We thank the Danish e-Infrastructure Consortium (DeiC), Martin Brandt, and Konrad Schindler for their support with computing resources.
@@ 0.A.1 Sampling strategy
ERA5. We selected 4 bands from the ERA5 dataset on GEE (average temperature, minimum temperature, maximum temperature and total precipitation). These statistics are monthly. Additionally, we computed the yearly average, minimum, maximum, and total, respectively, for each of these statistics. We collected 3 sets of statistics for the same month, previous month, and full year corresponding to the Sentinel-2 observation date ($3\times 4\ \text{bands}=12\ \text{bands}$).
Geolocation. We use a cyclic encoding of the latitude $\phi$ and longitude $\lambda$ corresponding to the center point as follows:
$\phi_{\sin}=\sin{(2\pi\phi/360)}$ (3)
$\phi_{\cos}=\cos{(2\pi\phi/360)}$ (4)
$\lambda_{\sin}=\sin{(2\pi\lambda/360)}$ (5)
$\lambda_{\cos}=\cos{(2\pi\lambda/360)}$ (6)
@@
$m_{\sin}=\sin{(2\pi m/12)}$ (7)
$m_{\cos}=\cos{(2\pi m/12)}$ (8)
Here $m$ corresponds to the month as an integer from 1 to 12.
@@ 0.A.2 Licenses of data sources
Appendix 0.B Implementation details
[…]
Semantic segmentation: We adopt a U-Net [39] architecture to evaluate the pretrained encoders. The upsampling block consists of a nearest neighbour upsampling followed by a convolutional layer with 3$\times$3 filter kernels, layer norm and a GELU activation layer.
Input bands for So2Sat20k: So2Sat20k only provides 10 instead of 12 bands. However, to make use of these additional multi-spectral bands, we filled in the missing bands (B1 and B9) with a copy of the band with the closest wavelength. I.e., we use a copy of B2 for B1, and B8A for B9. This could lead to a lower linear probing performance but should not affect results much when fine-tuning the entire encoder.
@@ Appendix 0.C Taster pretraining datasets: MMEarth100k and MMEarth64
To facilitate research in self-supervised learning with limited compute resources, we provide two subsets of the MMEarth dataset: MMEarth100k and MMEarth64. The MMEarth100k is a random subset of 100k locations with the full 128$\times$128 pixels and has a data volume of roughly 48GB. The MMEarth64 contains all 1.2M locations, but all modalities are center cropped to obtain rasters of 64$\times$64 pixels, which reduces the data volume roughly by factor 4 from 597GB to 152GB.
We provide results using these taster subsets to pretrain our MP-MAE approach with the same setting as used for pretraining on the full MMEarth dataset in Table A2. The only change required for MMEarth64 is the patch size. As the image size is halved from 112 to 56 pixels, we also half the patch size from 16 to 8 pixels which keeps the number of patches the same. For the downstream tasks that we consider in this evaluation, we observe that pretraining on MMEarth64 performs comparably to MMEarth, and even slightly better in some tasks. However, there might be downstream tasks that could profit from larger pretraining image sizes. Also alternative self-supervised pretraining strategies might benefit from having access to the larger input sizes. Pretraining on MMEarth100k leads to a consistent drop in performance, but still outperforms the ImageNet pretrained models on all tasks except the Cashew1k. From these initial results, we conclude that the diversity and number of samples are more important pretraining characteristics than the size of the input images for learning generalizing representations with our proposed MP-MAE.
Table A2: MP-MAE results on MMEarth taster datasets.
[…]
@@ 0.D.1.1 Analysis of the estimated task uncertainty
Table A3: Loss aggregation strategy. We evaluate the effect of the two loss aggregation strategies: Equal weighting vs. task uncertainty weighting when pretraining on MMEarth64.
[…]
0.D.2 Qualitative reconstruction results